jux has double times balls as joe and red has 3 times asjux's.altogether they have a total of 18 balls.whatis the algebraic equation?
Answer by stanbon(75887)

jux has double times balls as joe and red has 3 times asjux's.altogether they have a total of 18 balls.whatis the algebraic equation?
-------------------------
Equations:
x = 2j
r = 3x
-----
x + j + r = 18
------
Substitute::
x + x/2 + 3x = 18
-----
2x + x + 6x = 36
-----
9x = 36
x = 4
j has x/2 = 2 balls
r has 3x = 12 balls
------------------
Cheers,
Stan H.
------------